# Modelling subsonic turbulence with SPH-EXA

### v1

@@ II-C Volume elements
where $X$ is an arbitrary function. In SPH-EXA we adopted the GVE described in the work of [18], which has shown to provide a better treatment of discontinuities, stronger resistance to tensile instability, better partition of unity, and more accurate gradient evaluation:
@@
where $\rho^{0}_{a}=\sum_{b}m_{b}W_{ab}$ is the standard SPH density calculation. With Eqs. 17 and 18 we can simply calculate the particle density as $\rho_{a}=m_{a}/V_{a}$.
II-D Lagrangian compatibility
One of the most common artifacts in SPH simulations is tensile instability, which can, in some cases, prevent mixing and the development of hydrodynamical instabilities. Several approaches have been developed to deal with such problems [27, 28, 25], all of which require a departure to some extent from the pure Lagrangian formulation. In SPH-EXA we adopted a novel implementation [18], which, to our knowledge, is currently the only adaptive method that sacrifices Lagrangian compatibility exclusively in regions of the fluid at (and adjacent to) contact interphases, improving the accuracy at those regions, in turn. Hence, we define a self-adaptive parameter that is individual to each particle-neighbor pair as follows:
$\sigma_{ab}\equiv\begin{cases}0&\text{if~{}~{}$At_{ab}\leq At_{min}$}\,,\\ R(At_{ab}-At_{min})&\text{if~{}~{}$At_{min}\leq At_{ab}\leq At_{max}$}\,,\\ 1&\text{if~{}~{}$At_{ab}\geq At_{max}$}\,,\end{cases}$ (19)
with $R=(At_{max}-At_{min})^{-1}$ and $At_{ab}$ is the pairwise Atwood number, defined as $At_{ab}=\left|\frac{\rho_{a}-\rho_{b}}{\rho_{a}+\rho_{b}}\right|$. We use the values $At_{max}=0.2$ and $At_{min}=0.1$, as suggested in [18]. In this implementation $0\leq\sigma\leq 1$ (see Eqs. 22 and 23). In particular, when $\sigma=0$ the resulting SPH equations are fully Lagrangian, while $\sigma=1$ leads to a version less Lagrangian compatible but more tensile resistant. The parameter $\sigma$ appears in the momentum and energy SPH equations and has the effect of crossing and uncrossing the particle and neighbor’s magnitudes.
II-E Interpolation kernel
We use the sinc family ($S_{n}$) [29, 30] as interpolation kernels,
$W_{ab}=S_{n}=\frac{K}{h_{a}^{d}}\begin{dcases}1&q_{a}=0\\ \left[\operatorname{sinc}\left(\frac{\pi}{2}q_{a}\right)\right]^{n}&q_{a}\leq 2\\ 0&q_{a}>2\\ \end{dcases}$ (20)
where $K$ is a normalization constant, $h$ is the smoothing length, $d$ is the spatial dimension, $q_{a}$ is the normalized distance between neighbor particles ($q_{a}=|\mathbf{x}_{b}-\mathbf{x}_{a}|/h_{a}$), and $n$ is a user-defined real value that makes the kernel more centrally peaked as it increases. Usually $93<n<$ 9. It is known [31] that a necessary condition for an SPH kernel to avoid pairing instability is that its Fourier transform is definite positive. As such, they proposed the use of Wendland functions [32] as SPH kernels. These have been adopted by many researchers for this reason. The other side of the coin is that Wendland kernels need a large number of neighbors to perform accurate interpolations, which leads to considerable computational efforts. A possible alternative is to use pairing-resistant kernels. Despite not having an always positive Fourier transform, they are constructed to have small negative Fourier regions and push them to large wavenumbers. This allows these kernels to use a lower number of neighbors than Wendland functions, while still suppressing pairing instability. Sinc kernels are an example of such kernels. The simulations reported in this work were calculated choosing $n=6$ in Eq. 20.
II-F SPH-EXA hydrodynamic equations
@@
$\begin{split}\frac{dv_{i,a}}{dt}=&-\sum_{b}m_{b}\times\\ &\left[\frac{X_{a}^{2-\sigma}X_{b}^{\sigma}P_{a}}{\Omega_{a}m_{a}^{2}~{}k_{a}}A_{i,ab}(h_{a})+\frac{X_{b}^{2-\sigma}X_{a}^{\sigma}P_{b}}{\Omega_{b}m_{b}^{2}~{}k_{b}}A_{i,ab}(h_{b})\right]\,,\end{split}$ (22)
@@
$\begin{split}\frac{du_{a}}{dt}=&\frac{X_{a}^{2-\sigma}P_{a}}{m_{a}^{2}\Omega_{a}k_{a}}\sum_{b}\sum^{d}_{i=1}m_{b}X_{b}^{\sigma}\left[(\mathbf{v}_{a}-\mathbf{v}_{b})\cdot\mathbf{A}_{ab}(h_{a})\right]\\ &+\left(\frac{du_{a}}{dt}\right)^{AV}\,.\end{split}$ (23)
@@
Here $P_{a}$ is the pressure, $X_{a}=m_{a}/\rho^{0}_{a}$ the selected function for the GVE, $k_{a}$ the normalization of the GVE ($k_{a}\equiv\sum_{b}X_{b}W_{ab}$ in Eq. 17), $A_{i,ab}(h_{a})$ the IA term, and $\Omega_{a}$ the grad-h term. The parameter $\sigma$ controls how close the implementation is to a pure Lagrangian scheme (as discussed above). The last term in the energy equation includes the contributions of artificial viscosity.
II-G Initial conditions
We have designed an embedded initial conditions (IC) generator in SPH-EXA. In all our tests, we use a template of a relaxed box with $50^{3}$ particles. This box has previously been relaxed to achieve a glass-like configuration [33] with constant density and periodic boundary conditions. Replicating this box in different directions allows us to create on-the-fly initial conditions with a user-defined number of particles and in a distributed manner, only limited by the memory of the system. In this way, ICs can be scaled up just by changing one input argument and without the need to store large files, ensuring reproducibility and simplifying the process of performing suites of tests at scale.
All simulations presented here were performed on a box of side $L=1$, periodic boundary conditions, initial density $\rho=1$, and ’quasi’-isothermal equation of state with an adiabatic index $\gamma=1.001$. The fluid is stirred using the Ornstein-Uhlenbeck process [34, 35] following a parabolic amplitude field $\sim 1-(k-k_{c})^{2}$, with $k=|\mathbf{k}|$, and $k_{c}=2$, which corresponds to a length $L/2$, and vanishing for $k\leq 1$ and $k\geq 3$. This approach is implemented to reduce the impact of the stirring on the energy cascade, facilitating a natural transfer of energy from larger to smaller scales. We also use a natural mixing of $1/3$ compressive and $2/3$ solenoidal contributions. All simulations start in a homogenous state and are stirred for a total of $t=10$, until they reach the stationary state [^1] All our times are normalized with the sound-crossing time, which in our setup is $t_{sc}=L/c=1$. with an RMS Mach number of $M=0.3$.
III Results
To evaluate the weight and significance of the different elements that contribute to our SPH implementation, we conducted identical simulations while progressively incorporating more of these elements. Figure 1 shows the power spectrum of the velocity field of simulations of $250^{3}$ particles with increasing complexity, compared to the S3 simulation of [9] with $256^{3}$ particles and the classic SPH formulation (i.e. gradients calculated directly as kernel derivatives, no GVEs, and no viscosity switches), labeled STD in our plot.
Figure 1: Comparison of the power spectra of the velocity field for $M=0.3$ with different elements incorporated into the SPH solver. The dashed line is the theoretical Kolmogorov power spectrum of $\sim k^{-5/3}$. STD stands for standard SPH and is taken from [9] (simulation S3 in their Fig. 6, with $256^{3}$ particles). All other simulations were performed with SPH-EXA and $250^{3}$ particles. IA stands for including the integral formalism to calculate derivatives. VAV stands for variable AV and implies the inclusion of AV switches. GVE is for generalized volume elements, and AVC for AV cleaning. All these elements have been described in Sec. II.
The inclusion of IA recovers the correct slope in the inertial range of the Kolmogorov cascade for smaller wavenumbers, considerably extending it compared to STD, but is not enough to affect the classical excessive slope of $\sim-4$ at large wavenumbers. At these small scales something else is needed. The inclusion of GVE and artificial viscosity switches (labeled as VAV in Fig. 1 provides a step in the right direction. With these elements, we start to see the bottleneck effect as the excess of energy at increasing wavenumber, commonly encountered close to the resolution limit in mesh-based simulations. Finally, the inclusion of the artificial viscosity cleaner (AVC in Fig. 1) provides the final push, extending the inertial range almost a decade with respect STD, showing the bottleneck effect clearly, and having a slope in the dissipative regime similar to those of non-Lagrangian codes (see Fig. 4).
Figure 2 presents the power spectrum of the velocity field for simulations performed with SPH-EXA and with different particle counts at the end of each simulation ($t=10$). It is clear that with increasing resolution, the inertial range over the Kolmogorov scale is extended to larger wavenumber, and more importantly, that the rate of enhancement with increasing number of particles is not marginal or small, but considerable and consistent with that obtained with modern mesh-based and moving-mesh methods. This is a novel result, which means that SPH can indeed resolve smaller scales with higher resolution, in opposition to the results found with older SPH implementations by [9]. Figure 3 shows a thin slice of the velocity field for increasing resolution with SPH-EXA, from $200^{3}$ up to $3000^{3}$, where it is evident how we can resolve a delicate and rich mixture of large- and small-scale features.
Figure 2: Comparison of the power spectra of the velocity field at $t=10$ for $M=0.3$ at increasing resolution. Each line corresponds to an increase of a factor 2 in resolution with respect to the previous curve.
Figure 3: Slice of the module of the velocity field with increasing resolution and calculated with SPH-EXA.
Now we can directly compare our results with other state-of-the-art codes that are not purely Lagrangian. Figure 4 displays a comparison of the power spectra calculated with AREPO, GIZMO, and SPH-EXA. As can be seen, SPH-EXA is able to reproduce the same inertial range. We can also see the bottleneck effect to a very similar extent in all three codes. As this artificial excess of energy happens at a very similar wavenumber for similar resolutions, it points to the fact that the transition from the inertial range to the dissipation range happens at the same scale for a given resolution and for all three codes. It is worth noting that in Fig. 4 we compare codes with a similar amount of fluid elements, which does not necessarily correspond to same resolution. In SPH codes, the spatial resolution is defined by the smoothing length, which covers significantly more than just one particle. At $256^{3}$ cells, the spatial resolution of AREPO roughly corresponds to $1/256\sim 3.9\times 10^{-3}$, which is in fact similar to our $400^{3}$ SPH simulation, which has an average smoothing length of $h\sim 3.6\times 10^{-3}$. This means that if we compare at the same resolution, and according to the results of Fig. 4, SPH-EXA captures a considerably larger inertial range of the Kolmogorov cascade than AREPO. Likewise, the simulation of AREPO with $512^{3}$ cells has an approximate resolution of $\sim 1.9\times 10^{-3}$, which corresponds to our $800^{3}$ simulation ($h\sim 1.8\times 10^{-3}$) represented with a dashed line in Fig. 4.
Figure 4: Comparison of the power spectra of the velocity field for $M=0.3$ for different codes. AREPO [9] and GIZMO [20] curves are from their respective papers.
Figure 5 shows, from top to bottom, the module of the velocity, density, and enstrophy of the fluid at the end of our simulations. The two columns on the left are an excerpt of the results with AREPO and the SPH used in [9]. It is clear from these two columns that the SPH used in that work was not able to resolve subsonic turbulence, since it is missing all small-scale features. The two columns on the right show the results of SPH-EXA with $250^{3}$ and $1000^{3}$ particles. The similarities with the results of AREPO are evident, and there is a rich mix of fine structures that becomes richer with increasing resolution. This can also be seen in Fig. 3, which illustrates a slice of the module of the velocity field at $t=10$ for different resolutions.
Figure 5: Comparison of SPH-EXA results with AREPO and SPH-tav in [9]. From top to bottom: module of velocity, density, and enstrophy. The two columns on the left are an excerpt of the original results of AREPO and their SPH with reduced artificial viscosity (both with $256^{3}$ fluid elements) of [9] shown in their Fig. 4. The two columns on the right are the results of SPH-EXA with $250^{3}$ and $1000^{3}$
Finally, Fig. 6 is a detailed image of a slice of the module of velocity of the $3000^{3}$ simulation where we can appreciate the very small-scale features resolved with SPH-EXA.
Figure 6: Slice of the module of the velocity field at $t=10$ for a simulation with $3000^{3}$ particles. Color follows the same scale as Fig. 3 but in grey.
IV Summary and conclusions
SPH has matured enough to be competitive in simulations of subsonic turbulence, obtaining results comparable to Eulerian codes. In particular, SPH-EXA combines a modern SPH formulation with the latest techniques in parallelization and the use of accelerators. This allowed us to overcome historical limitations of SPH, such as excessive dissipation and inaccurate gradient evaluations, which have traditionally prevented SPH from matching the performance of grid-based and moving-mesh codes in this field. Our findings not only align well with results from non-Lagrangian codes like AREPO and GIZMO, but also illustrate the robustness of SPH in handling complex turbulent flows at resolutions previously thought unattainable with this method.
We demonstrated that incorporating various components that address problems at multiple simulation scales collectively achieves these results. Isolated enhancements in the SPH framework cannot tackle the subsonic turbulence; however, a synergy of these improvements can. In particular, our combination of an integral approach to derivatives, generalized volume elements, artificial viscosity switches, and artificial viscosity cleaner shows excellent results.
Finally, we proved that the inertial range of the power spectra significantly spans over larger vawenumbers with increasing resolution, exhibiting the same features as other non-Lagrangian state-of-the-art codes. These improvements have allowed SPH to finally tap into its inherent advantages, such as its intrinsic Lagrangian nature, excellent conservation properties, and adaptability to varied physical scenarios, thus paving the way for its application in more demanding astrophysical contexts.
Acknowledgment
This work has been supported by the Swiss Platform for Advanced Scientific Computing (PASC) project ”SPH-EXA: Optimizing Smoothed Particle Hydrodynamics for Exascale Computing”. It has also been carried out as part of the SKACH consortium through funding from SERI and supported by the Spanish MINECO grant PID2020-117252GB-100 and by the AGAUR/Generalitat de Catalunya grant SGR-386/2021. We also acknowledge the financial support of the European Research Council through the ERC Synergy Grant “ECOGAL” (project ID 855130), from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) “STRUCTURES”, and from the German Ministry for Economic Affairs and Climate Action in project “MAINN” (funding ID 50OO2206). The numerical calculations have been supported by the EuroHPC-JU Extreme Scale Access Mode through the project ”TGSF: The role of turbulence and gravity in star formation: unveiling the sonic scale with Smoothed Particle Hydrodynamics” (proposal ID: EHPC-EXT-2023E01-031). The authors acknowledge the support of the LUMI Supercomputer Center, the Swiss National Supercomputing Center (CSCS - allocation c32), and the Center for Scientific Computing (sciCORE) at the University of Basel, where these calculations were performed.

### v2

@@ A.3 Volume elements
where $X$ is an arbitrary function. In SPH-EXA, we adopted the GVE described in the work of García-Senz et al. (2022), which has been shown to provide a better treatment of discontinuities, stronger resistance to tensile instability, better partition of unity, and more accurate gradient evaluation:
@@
where $\rho^{0}_{a}=\sum_{b}m_{b}W_{ab}$ is the standard SPH density calculation. With Eqs. 19 and 20 we can simply calculate the particle density as $\rho_{a}=m_{a}/V_{a}$.
A.4 SPH-EXA hydrodynamic equations
@@
[…]
@@
$\begin{split}\frac{du_{a}}{dt}=&\frac{X_{a}^{2-\sigma}P_{a}}{m_{a}^{2}\Omega_{a}k_{a}}\sum_{b}\sum^{d}_{i=1}m_{b}X_{b}^{\sigma}\left[(\mathbf{v}_{a}-\mathbf{v}_{b})\cdot\mathbf{A}_{ab}(h_{a})\right]+\left(\frac{du_{a}}{dt}\right)^{AV}\,.\end{split}$ (23)
@@
Here $P_{a}$ is the pressure, $X_{a}=m_{a}/\rho^{0}_{a}$ the selected function for the GVE, $k_{a}$ the normalization of the GVE ($k_{a}\equiv\sum_{b}X_{b}W_{ab}$ in Eq. 19), $A_{i,ab}(h_{a})$ the IA term, and $\Omega_{a}$ the grad-h term. The parameter $\sigma$ controls how close the implementation is to a pure Lagrangian scheme (see García-Senz et al. 2022). The last term in the energy equation includes the contributions of artificial viscosity.
Appendix B The relevance of accurate grad-h terms
Figure 7: Effect of the accuracy in grad-h terms. The top row shows the value of the grad-h correction $\Omega$ for all particles in a $250^{3}$ simulation at $t=10$. Left, $\Omega$ is calculated using the density field directly. Right, using the differences in the density field. The red lines show the mean value of $\Omega$. The bottom row shows a thin slice of the resulting density field.
Grad-h terms are an important addition to the SPH technique when using variable smoothing lengths ($h$) (Springel and Hernquist 2002; Monaghan 2002). It takes into account the spatial variation of $h$ as a corrective term in the momentum (Eq. 22) and energy (Eq. 23) equations:
$\Omega_{a}=\left[1+\frac{h_{a}}{3\rho_{a}}\frac{\partial\rho_{a}}{\partial h_{a}}\right]^{-1}\,.$ (24)
These corrections are a standard method to improve accuracy and consistency in particle interpolation, reducing numerical instability in variable-density simulations, and as such, they are included by default in most astrophysics SPH codes.
During the subsonic simulations presented above, we noticed that the density field appeared ”granulated” only when grad-h corrections were used jointly with SLR dissipation (see Fig. 7, left column). Initially, we suspected that the new addition (the SLR dissipation control) was the source of the problem, but it turned out that the problematic term was how the grad-h correction was calculated. After inspecting the values of this correction we noticed that the average was systematically below 1 and with a relatively large spread. For example for the $250^{3}$ simulation, $\left<\Omega\right>=0.98\pm 0.16$. This doesn’t seem much, but it is exactly the kind of ”small” inconsistency that becomes relevant in subsonic turbulence, in particular when the dissipation formulation is designed to stop damping that noise.
In low-Mach turbulence, physical density fluctuations are weak, so any residual inconsistency in the pressure–density–$h$ coupling becomes visible. We can write the grad-h field as:
$\Omega_{a}=\bar{\Omega}+\delta\Omega_{a}\,.$ (25)
Then, expanding the multiplicative term in the SPH equations:
$\frac{1}{\Omega_{a}}=\frac{1}{\bar{\Omega}}\left(1-\frac{\delta\Omega_{a}}{\bar{\Omega}}+\mathcal{O}(\delta\Omega_{a}^{2})\right)$ (26)
The $1/\Omega_{a}$ multiplicative factor that scales the pressure forces has two error sources: a biased mean and a noisy scatter. The first is just a constant scaling affecting the effective compressibility and shifting the global balance slightly. It shows as a systematic drift: slightly different RMS Mach, different effective forcing response, and slightly different spectra cutoff location, but doesn’t introduce particle-scale structure by itself. The second is the actual source of the granularity, since it is a multiplicative noise field coupled to the pressure acceleration. Since $\delta\Omega_{a}$ varies on the inter-particle scale and correlates with local disorder, it acts like a high-$k$ forcing of the compressive jitter:
$\mathbf{a}_{a}^{P}\approx\frac{1}{\bar{\Omega}}\mathbf{a}^{P}_{\Omega=1}-\frac{1}{\bar{\Omega}^{2}}\delta\Omega_{a}\mathbf{a}^{P}_{\Omega=1}+\dots$ (27)
If these corrections are in fact a noisy field, we get a noisy acceleration and particle jitter. This leads to noisy density estimates, noisy smoothing length, and back to noisy grad-h corrections, creating a feedback loop that distorts the field completely. This is usually masked by baseline AV, but becomes apparent when SLR suppresses dissipation in smooth regions.
According to Eq. 24 we need to evaluate $\partial\rho_{a}/\partial h_{a}$ and, as far as we know, the literature always implements it as a direct derivative of the density field. This is well known to have considerable E0 errors, and a typical trick is to include a scalar in the derivative to calculate it using the differences of the field, rather than the field directly. This has been used extensively to yield built-in conservation in SPH and zero derivatives when the field is constant. For example:
$\frac{\partial\rho}{\partial h}=\frac{1}{\phi}\left[\frac{\partial(\phi\rho)}{\partial h}-\rho\frac{\partial\phi}{\partial h}\right]\,$, (28)
where $\phi$ could be any other field or scalar. If we simply choose $\phi=1$:
$\frac{\partial\rho_{a}}{\partial h_{a}}=\sum_{b}V_{b}\rho_{b}\frac{\partial W_{ab}(h_{a})}{\partial h_{a}}-\rho_{a}\sum_{b}V_{b}\frac{\partial W_{ab}(h_{a})}{\partial h_{a}}\,.$ (29)
The first term on the RHS of Eq. 29 is the common grad-h term that is used to calculate $\Omega_{a}$. The second term in the RHS of Eq. 29 is the additional correction to evaluate grad-h as differences and reduce its noise.
The practical implementation of this term depends on which volume elements are used. In SPH-EXA, we corrected the calculation of $\partial\rho_{a}/\partial h_{a}$, amending formula A17 from García-Senz et al. (2022). Defining:
$\Psi_{a}\equiv\rho_{a}\sum_{b}V_{b}\frac{\partial W_{ab}(h_{a})}{\partial h_{a}}\,$, (30)
we obtain:
$\begin{split}\frac{\partial\rho_{a}}{\partial h_{a}}=\left[\frac{\rho_{a}}{\rho^{0}_{a}}-X_{a}W_{aa}(h_{a})\right]&\sum_{b}m_{b}\frac{\partial W_{ab}(h_{a})}{\partial h_{a}}-\rho^{0}_{a}\Psi_{a}+\\ \frac{m_{a}}{X_{a}}&\sum_{b}X_{b}\frac{\partial W_{ab}(h_{a})}{\partial h_{a}}-\rho_{a}\Psi_{a}\,.\end{split}$ (31)
Note that if this formulation is used, self-contributions of particle $a$ must not be included. There are different ways to implement this correction, as well as different possibilities for the field $\phi$. We opted for the easiest and least invasive methodology.
As a result, Fig. 7 (right column) shows that the average value of $\Omega$ improved considerably, becoming $\left<\Omega\right>=1.00\pm 0.04$, and the density field was recovered.
Appendix C Second-order velocity structure function
Figure 8: Mass-weighted turbulent Mach number as a function of spatial scale for a turbulent box simulation with a resolution of $2000^{3}$ particles. The dotted line indicates the RMS velocity of the simulation. The dashed line is a power-law fit to the turbulent cascade region.
The SPH method is particularly advantageous for a wide range of mass-weighted analysis methods. Most importantly, it removes the need for special tracer particles as the SPH particles themselves serve as tracers. This allows for an easy computation of several quantities like the finite-time Lyapunov exponent and mass-weighted structure functions. The structure function is defined as
${SF}_{2}(d/L)=\left\langle|\@vec{v}(\@vec{r})-\@vec{v}(\@vec{r}+\@vec{d}/L)|^{2}\right\rangle_{\@vec{r}}$ (32)
where d=|d(r)|d=|\@vec{d}(\@vec{r})| is the distance between the two particles in a pair, and we compute the average over many SPH particle pairs, which makes this a mass-weighted structure function. $L$ is the turbulent driving scale. From this, one can compute the turbulent Mach number as
$M_{t}(d/L)=\sqrt{{SF}_{2}(d/L)/(2c_{s}^{2})}$ (33)
As an example, Figure 8 shows the turbulent Mach number as a function of the spatial scale for one of our SPH-EXA simulations. The computation was done using $10^{9}$ particle pairs per snapshot and was averaged over $41$ snapshots of the simulation, which is sufficient to achieve statistical convergence.
In the regime of the turbulent cascade, we find a power law with an exponent of $0.333$. This perfectly matches the expected result of $1/3$ for subsonic Kolmogorov turbulence. Together with the density PDF shown in Figure 4, this demonstrates that SPH-EXA is capable of producing accurate, mass-weighted statistics for subsonic turbulence in a much more natural way than comparable mesh codes.
Appendix D High-resolution view of the velocity field
Figures 9 to 11 show a detailed image of a slice of the velocity magnitude, density, and enstrophy for the $3200^{3}$ simulation, respectively, where we can appreciate the very small-scale features resolved with SPH-EXA. All slices have a thickness of $2h$ and they are centered on $z=0$. Particles are displayed directly (no particle-to-grid interpolation is applied) and each slice contains $\sim 5.84\times 10^{7}$ particles.
Figure 9: Velocity magnitude $|\mathbf{v}|$ at $t=10$ for the $3200^{3}$ simulation, shown in a slice of thickness $2h$ centered on $z=0$.
Figure 10: Density at $t=10$ for the $3200^{3}$ simulation, shown in a slice of thickness $2h$ centered on $z=0$.
Figure 11: Enstrophy at $t=10$ for the $3200^{3}$ simulation, shown in a slice of thickness $2h$ centered on $z=0$.